# On the Mode-Seeking Properties of Langevin Dynamics

### v2

@@ A.1 Proof of Theorem 1: Langevin Dynamics under Gaussian Mixtures
Without loss of generality, we assume that ${\bm{\mu}}_{0}=\bm{0}_{d}$ for simplicity. Let $r$ and $n$ respectively denote the rank and nullity of the vector space $\left\{{{\bm{\mu}}_{i}}\right\}_{i\in[k]}$, then we have $r+n=d$ and $0\leq r\leq k=o(d)$. Denote ${\mathbf{R}}\in{\mathbb{R}}^{d\times r}$ an orthonormal basis of the vector space $\left\{{{\bm{\mu}}_{i}}\right\}_{i\in[k]}$, and denote ${\mathbf{N}}\in{\mathbb{R}}^{d\times n}$ an orthonormal basis of the null space of $\left\{{{\bm{\mu}}_{i}}\right\}_{i\in[k]}$. Now consider decomposing the sample ${\mathbf{x}}_{t}$ by
@@
It is worth noting that since ${\mathbf{N}}^{T}{\bm{\mu}}_{i}=\bm{0}_{n}$. To show $\left\|{{\mathbf{x}}_{t}-{\bm{\mu}}_{i}}\right\|^{2}>\frac{{\nu}_{0}^{2}+{\nu}_{\max}^{2}}{2}d$, it suffices to prove
$\left\|{\mathbf{n}_{t}}\right\|^{2}>\frac{{\nu}_{0}^{2}+{\nu}_{\max}^{2}}{2}d.$
We start by proving that the initialization of the state ${\mathbf{x}}_{0}$ has a large norm on the null space with high probability in the following proposition. Throughout the proof, the notation $\Omega(d)$ refers to $\Omega(d)\geq cd$, for the constant $c$ defined as
$c=\min\left\{\frac{1}{2}\left(\frac{\nu_{0}^{2}-\nu_{\max}^{2}}{8\nu_{0}^{2}}\right)^{2},\frac{1}{8}\left(\log\left(\frac{\nu_{\max}^{2}}{\nu_{0}^{2}}\right)-\frac{\nu_{\max}^{2}}{2\nu_{0}^{2}}+\frac{\nu_{0}^{2}}{2\nu_{\max}^{2}}\right),\frac{1}{32},\frac{(\nu_{0}^{2}-\nu_{\max}^{2})^{2}}{32\nu_{0}^{2}(\nu_{0}^{2}+\nu_{\max}^{2})}\right\},$ (6)
when $d$ satisfies
$d\geq\max\left\{8\left(\log\left(\frac{\nu_{\max}^{2}}{\nu_{0}^{2}}\right)-\frac{\nu_{\max}^{2}}{2\nu_{0}^{2}}+\frac{\nu_{0}^{2}}{2\nu_{\max}^{2}}\right)^{-1}\log\left(\frac{3\nu_{0}^{3}}{w_{0}\min_{i\in[k]}\nu_{i}^{2}}\right),\frac{8\nu_{0}^{2}(3\nu_{0}^{2}+\nu_{\max}^{2})}{\pi(\nu_{0}^{2}-\nu_{\max}^{2})^{2}}\right\}.$ (7)
Proposition 2.
Suppose that a sample ${\mathbf{x}}_{0}$ is initialized in the distribution $P^{(0)}$, i.e., ${\mathbf{x}}_{0}\sim P^{(0)}$, then for any constant ${\nu}_{\max}<{\nu}_{0}$, with probability at least $1-\exp(-\Omega(d))$, we have $\left\|{\mathbf{n}_{0}}\right\|^{2}\geq\frac{3{\nu}_{0}^{2}+{\nu}_{\max}^{2}}{4}d$.
Proof of Proposition 2.
Since ${\mathbf{x}}_{0}\sim P^{(0)}={\mathcal{N}}(\bm{0}_{d},{\nu}_{0}^{2}{\bm{I}}_{d})$ and ${\mathbf{N}}^{T}{\mathbf{N}}={\bm{I}}_{n}$, we know $\mathbf{n}_{0}={\mathbf{N}}^{T}{\mathbf{x}}_{0}\sim{\mathcal{N}}(\bm{0}_{n},{\nu}_{0}^{2}{\bm{I}}_{n})$. Therefore, by Lemma 1 we can bound
${\mathbb{P}}\left(\left\|{\mathbf{n}_{0}}\right\|^{2}\leq\frac{3{\nu}_{0}^{2}+{\nu}_{\max}^{2}}{4}d\right)$ $={\mathbb{P}}\left(\frac{\left\|{\mathbf{n}_{0}}\right\|^{2}}{{\nu}_{0}^{2}}\leq d-2\sqrt{d\cdot\left(\frac{{\nu}_{0}^{2}-{\nu}_{\max}^{2}}{8{\nu}_{0}^{2}}\right)^{2}d}\right)$
$\leq{\mathbb{P}}\left(\frac{\left\|{\mathbf{n}_{0}}\right\|^{2}}{{\nu}_{0}^{2}}\leq n-2\sqrt{n\left(\frac{{\nu}_{0}^{2}-{\nu}_{\max}^{2}}{8{\nu}_{0}^{2}}\right)^{2}\frac{d}{2}}\right)$
$\leq\exp\left(-\left(\frac{{\nu}_{0}^{2}-{\nu}_{\max}^{2}}{8{\nu}_{0}^{2}}\right)^{2}\frac{d}{2}\right),$
where the second last step follows from the assumption $d-n=r=o(d)$. Hence we complete the proof of Proposition 2. ∎
Then, with the assumption that the initialization satisfies $\left\|{\mathbf{n}_{0}}\right\|^{2}\geq\frac{3{\nu}_{0}^{2}+{\nu}_{\max}^{2}}{4}d$, the following proposition shows that $\left\|{\mathbf{n}_{t}}\right\|$ remains large with high probability.
Proposition 3.
Consider a data distribution $P$ satisfies the constraints specified in Theorem 1. We follow the Langevin dynamics for $T=\exp({\mathcal{O}}(d))$ steps. Suppose that the initial sample satisfies $\left\|{\mathbf{n}_{0}}\right\|^{2}\geq\frac{3{\nu}_{0}^{2}+{\nu}_{\max}^{2}}{4}d$, then with probability at least $1-T\cdot\exp(-\Omega(d))$, we have that $\left\|{\mathbf{n}_{t}}\right\|^{2}>\frac{{\nu}_{0}^{2}+{\nu}_{\max}^{2}}{2}d$ for all $t\in\{0\}\cup[T]$.
Proof of Proposition 3.
@@
If $\delta_{t}>{\nu}_{0}^{2}$, with probability at least $1-\exp(-\Omega(d))$, for $\mathbf{n}_{t}$ satisfying equation 5, we have $\left\|{\mathbf{n}_{t}}\right\|^{2}\geq\frac{3{\nu}_{0}^{2}+{\nu}_{\max}^{2}}{4}d$ regardless of the previous state ${\mathbf{x}}_{t-1}$.
@@
$\leq\exp\left(-\left(\frac{{\nu}_{0}^{2}-{\nu}_{\max}^{2}}{8{\nu}_{0}^{2}}\right)^{2}\frac{d}{2}\right),$
where the second last step follows from the assumption $d-n=r=o(d)$. Hence we complete the proof of Lemma 3. ∎
We then consider the case when $\delta_{t}\leq{\nu}_{0}^{2}$. Let $\mathbf{r}:={\mathbf{R}}^{T}{\mathbf{x}}$ and $\mathbf{n}:={\mathbf{N}}^{T}{\mathbf{x}}$, then ${\mathbf{x}}={\mathbf{R}}\mathbf{r}+{\mathbf{N}}\mathbf{n}$. We first show that when $\left\|{\mathbf{n}}\right\|^{2}\geq\frac{{\nu}_{0}^{2}+{\nu}_{\max}^{2}}{2}d$, $P^{(i)}({\mathbf{x}})$ is exponentially smaller than $P^{(0)}({\mathbf{x}})$ for all $i\in[k]$ in the following lemma.
@@
Given that $\left\|{\mathbf{n}}\right\|^{2}\geq\frac{{\nu}_{0}^{2}+{\nu}_{\max}^{2}}{2}d$ and $\left\|{{\bm{\mu}}_{i}}\right\|^{2}\leq\frac{{\nu}_{0}^{2}-{\nu}_{i}^{2}}{2}\left(\log\left(\frac{{\nu}_{i}^{2}}{{\nu}_{0}^{2}}\right)-\frac{{\nu}_{i}^{2}}{2{\nu}_{0}^{2}}+\frac{{\nu}_{0}^{2}}{2{\nu}_{i}^{2}}\right)d$ for all $i\in[k]$, we have $\frac{P^{(i)}({\mathbf{x}})}{P^{(0)}({\mathbf{x}})}\leq\exp(-\Omega(d))$ for all $i\in[k]$.
@@
Hence, for $\left\|{\mathbf{n}}\right\|^{2}\geq\frac{{\nu}_{0}^{2}+{\nu}_{\max}^{2}}{2}d$ and $\left\|{{\bm{\mu}}_{i}}\right\|^{2}\leq\frac{{\nu}_{0}^{2}-{\nu}_{i}^{2}}{2}\left(\log\left(\frac{{\nu}_{i}^{2}}{{\nu}_{0}^{2}}\right)-\frac{{\nu}_{i}^{2}}{2{\nu}_{0}^{2}}+\frac{{\nu}_{0}^{2}}{2{\nu}_{i}^{2}}\right)d$, we have
@@
$\leq\exp\left(-\left(\log\left(\frac{{\nu}_{i}^{2}}{{\nu}_{0}^{2}}\right)-\frac{{\nu}_{i}^{2}}{2{\nu}_{0}^{2}}+\frac{{\nu}_{0}^{2}}{2{\nu}_{i}^{2}}\right)\frac{d}{4}\right).$
Notice that for function $f(z)=\log z-\frac{z}{2}+\frac{1}{2z}$, we have $f(1)=0$ and $\frac{{\,\mathrm{d}}}{{\,\mathrm{d}}z}f(z)=\frac{1}{z}-\frac{1}{2}-\frac{1}{2z^{2}}=-\frac{1}{2}\left(\frac{1}{z}-1\right)^{2}<0$ when $z\in(0,1)$. Thus, $\log\left(\frac{{\nu}_{i}^{2}}{{\nu}_{0}^{2}}\right)-\frac{{\nu}_{i}^{2}}{2{\nu}_{0}^{2}}+\frac{{\nu}_{0}^{2}}{2{\nu}_{i}^{2}}$ is a positive constant for ${\nu}_{i}<{\nu}_{0}$, i.e., $\rho_{i}({\mathbf{x}})=\exp(-\Omega(d))$. Therefore we finish the proof of Lemma 4. ∎
@@
Suppose $\delta_{t}\leq{\nu}_{0}^{2}$ and $\left\|{\mathbf{n}_{t-1}}\right\|^{2}>36{\nu}_{0}^{2}d$, then for $\mathbf{n}_{t}$ following from equation 5, we have $\left\|{\mathbf{n}_{t}}\right\|^{2}\geq{\nu}_{0}^{2}d$ with probability at least $1-\exp(-\Omega(d))$.
@@
$=\mathbf{n}_{t-1}-\frac{\delta_{t}}{2}\sum_{i=0}^{k}\frac{P^{(i)}({\mathbf{x}}_{t-1})}{P({\mathbf{x}}_{t-1})}\cdot\frac{{\mathbf{N}}^{T}({\mathbf{x}}_{t-1}-{\bm{\mu}}_{i})}{{\nu}_{i}^{2}}+\sqrt{\delta_{t}}\bm{\epsilon}_{t}^{(\mathbf{n})}$
$=\left(1-\frac{\delta_{t}}{2}\sum_{i=0}^{k}\frac{P^{(i)}({\mathbf{x}}_{t-1})}{P({\mathbf{x}}_{t-1})}\cdot\frac{1}{{\nu}_{i}^{2}}\right)\mathbf{n}_{t-1}+\sqrt{\delta_{t}}\bm{\epsilon}_{t}^{(\mathbf{n})}.$ (8)
By Lemma 4, we have $\frac{P^{(i)}({\mathbf{x}}_{j-1})}{P^{(0)}({\mathbf{x}}_{j-1})}\leq\exp(-\Omega(d))$ for all $i\in[k]$, therefore
$1-\frac{\delta_{t}}{2}\sum_{i=0}^{k}\frac{P^{(i)}({\mathbf{x}}_{t-1})}{P({\mathbf{x}}_{t-1})}\cdot\frac{1}{{\nu}_{i}^{2}}\geq 1-\frac{\delta_{t}}{2}\cdot\frac{1}{{\nu}_{0}^{2}}-\frac{\delta_{t}}{2}\sum_{i\in[k]}\frac{w_{i}P^{(i)}({\mathbf{x}}_{t-1})}{w_{0}P^{(0)}({\mathbf{x}}_{t-1})}\cdot\frac{1}{{\nu}_{i}^{2}}\geq 1-\frac{1}{2}-\exp(-\Omega(d))>\frac{1}{3}.$ (9)
@@
For any $t\geq 1$, given that $\delta_{j}\leq{\nu}_{0}^{2}$ and $\frac{{\nu}_{0}^{2}+{\nu}_{\max}^{2}}{2}d\leq\left\|{\mathbf{n}_{j-1}}\right\|^{2}\leq 36{\nu}_{0}^{2}d$ for all $j\in[t]$ and $\left\|{{\bm{\mu}}_{i}}\right\|^{2}\leq\frac{{\nu}_{0}^{2}-{\nu}_{i}^{2}}{2}\left(\log\left(\frac{{\nu}_{i}^{2}}{{\nu}_{0}^{2}}\right)-\frac{{\nu}_{i}^{2}}{2{\nu}_{0}^{2}}+\frac{{\nu}_{0}^{2}}{2{\nu}_{i}^{2}}\right)d$ for all $i\in[k]$, we have $\left\|{\hat{\mathbf{n}}_{t}-\mathbf{n}_{t}}\right\|\leq\frac{t}{\exp(\Omega(d))}\sqrt{d}$.
@@
$=\left\|{\left(1-\frac{\delta_{j}}{2{\nu}_{0}^{2}}\right)(\hat{\mathbf{n}}_{j-1}-\mathbf{n}_{j-1})+\frac{\delta_{j}}{2}\sum_{i\in[k]}\frac{P^{(i)}({\mathbf{x}}_{j-1})}{P({\mathbf{x}}_{j-1})}\left(\frac{1}{{\nu}_{i}^{2}}-\frac{1}{{\nu}_{0}^{2}}\right)\mathbf{n}_{j-1}}\right\|$
$\leq\left(1-\frac{\delta_{j}}{2{\nu}_{0}^{2}}\right)\left\|{\hat{\mathbf{n}}_{j-1}-\mathbf{n}_{j-1}}\right\|+\sum_{i\in[k]}\frac{\delta_{j}}{2}\frac{P^{(i)}({\mathbf{x}}_{j-1})}{P({\mathbf{x}}_{j-1})}\left(\frac{1}{{\nu}_{i}^{2}}-\frac{1}{{\nu}_{0}^{2}}\right)\left\|{\mathbf{n}_{j-1}}\right\|$
$\leq\left\|{\hat{\mathbf{n}}_{j-1}-\mathbf{n}_{j-1}}\right\|+\sum_{i\in[k]}\frac{\delta_{j}}{2}\frac{P^{(i)}({\mathbf{x}}_{j-1})}{P^{(0)}({\mathbf{x}}_{j-1})}\left(\frac{1}{{\nu}_{i}^{2}}-\frac{1}{{\nu}_{0}^{2}}\right)6{\nu}_{0}\sqrt{d}.$
By Lemma 4, we have $\frac{P^{(i)}({\mathbf{x}}_{j-1})}{P^{(0)}({\mathbf{x}}_{j-1})}\leq\exp(-\Omega(d))$ for all $i\in[k]$, hence we obtain a recursive bound
@@
Given that $\left\|{\hat{\mathbf{n}}_{0}}\right\|^{2}\geq\frac{3{\nu}_{0}^{2}+{\nu}_{\max}^{2}}{4}d$, we have $\left\|{\hat{\mathbf{n}}_{t}}\right\|^{2}\geq\frac{5{\nu}_{0}^{2}+3{\nu}_{\max}^{2}}{8}d$ with probability at least $1-\exp\left(-\Omega(d)\right)$.
@@
$\leq{\mathbb{P}}\left(\left\|{\bm{\epsilon}}\right\|^{2}\leq(n-1)-2\sqrt{(n-1)\left(\frac{{\nu}_{0}^{2}-{\nu}_{\max}^{2}}{8{\nu}_{0}^{2}}\right)^{2}\frac{d}{2}}\right)$
$\leq\exp\left(-\left(\frac{{\nu}_{0}^{2}-{\nu}_{\max}^{2}}{8{\nu}_{0}^{2}}\right)^{2}\frac{d}{2}\right),$
@@
Upon having all the above lemmas, we are now ready to establish Proposition 3 by induction. Suppose the theorem holds for all $T$ values of $11,\cdots,T-$ 1. We consider the following 3 cases:
  • If there exists some $t\in[T]$ such that $\delta_{t}>{\nu}_{0}^{2}$, by Lemma 3 we know that with probability at least $1-\exp(-\Omega(d))$, we have $\left\|{\mathbf{n}_{t}}\right\|^{2}\geq\frac{3{\nu}_{0}^{2}+{\nu}_{\max}^{2}}{4}d$, thus the problem reduces to the two sub-arrays $\mathbf{n}_{0},\cdots,\mathbf{n}_{t-1}$ and $\mathbf{n}_{t},\cdots,\mathbf{n}_{T}$, which can be solved by induction.
  • Suppose $\delta_{t}\leq{\nu}_{0}^{2}$ for all $t\in[T]$. If there exists some $t\in[T]$ such that $\left\|{\mathbf{n}_{t-1}}\right\|^{2}>36{\nu}_{0}^{2}d$, by Lemma 5 we know that with probability at least $1-\exp(-\Omega(d))$, we have $\left\|{\mathbf{n}_{t}}\right\|^{2}\geq{\nu}_{0}^{2}d>\frac{3{\nu}_{0}^{2}+{\nu}_{\max}^{2}}{4}d$, thus the problem similarly reduces to the two sub-arrays $\mathbf{n}_{0},\cdots,\mathbf{n}_{t-1}$ and $\mathbf{n}_{t},\cdots,\mathbf{n}_{T}$, which can be solved by induction.
@@
    By Lemma 8 we have that with probability at least $1-\exp(-\Omega(d))$,

### v3

@@ A.2 Proof of Theorem 2
The proof of Theorem 2 follows from a similar framework to the proof of Theorem 1. Let $r$ and $n$ respectively denote the rank and nullity of the vector space $\left\{{{\bm{\mu}}_{i}}\right\}_{i\in[k]}$, then we have $r+n=d$ and $0\leq r\leq k=o(d)$. Denote ${\mathbf{R}}\in{\mathbb{R}}^{d\times r}$ an orthonormal basis of the vector space $\left\{{{\bm{\mu}}_{i}}\right\}_{i\in[k]}$, and denote ${\mathbf{N}}\in{\mathbb{R}}^{d\times n}$ an orthonormal basis of the null space of $\left\{{{\bm{\mu}}_{i}}\right\}_{i\in[k]}$. Now consider decomposing the sample ${\mathbf{x}}_{t}$ by
@@
By Definition 2, since $\mathbf{n}_{t}$ is the projection onto the null space of $\left\{{{\bm{\mu}}_{i}}\right\}_{i\in[k]}$, we have
$\left\|{{\mathbf{x}}_{t}}\right\|_{\left\{{{\bm{\mu}}_{i}}\right\}_{i\in[k]}}=\min_{\lambda_{1},\cdots,\lambda_{k}}\left\|{{\mathbf{x}}_{t}-\sum_{i=1}^{k}\lambda_{i}{\bm{\mu}}_{i}}\right\|=\left\|{\mathbf{n}_{t}}\right\|.$
Then, with the assumption that the initialization satisfies $\left\|{\mathbf{n}_{0}}\right\|^{2}\geq\frac{3{\nu}_{0}^{2}+{\nu}_{\max}^{2}}{4}d$, the objective is to show that $\left\|{\mathbf{n}_{t}}\right\|$ remains large with high probability.
@@
If $\delta_{t}>{\nu}_{0}^{2}$, with probability at least $1-\exp(-0.04d)$, for $\mathbf{n}_{t}$ satisfying equation 13, we have $\left\|{\mathbf{n}_{t}}\right\|^{2}\geq\frac{3{\nu}_{0}^{2}+{\nu}_{\max}^{2}}{4}d$ regardless of the previous state ${\mathbf{x}}_{t-1}$.
@@
$\leq\exp\left(-\left(\frac{{\nu}_{0}^{2}-{\nu}_{\max}^{2}}{8{\nu}_{0}^{2}}\right)^{2}\frac{d}{2}\right)\leq\exp\left(-\frac{d}{24}\right),$
Hence we complete the proof of Lemma 9. ∎
We then consider the case when $\delta_{t}\leq{\nu}_{0}^{2}$. We first show that when $\left\|{\mathbf{n}}\right\|^{2}\geq\frac{{\nu}_{0}^{2}+{\nu}_{\max}^{2}}{2}d$, $P^{(i)}({\mathbf{x}})$ is exponentially smaller than $P^{(0)}({\mathbf{x}})$ for all $i\in[k]$ in the following lemma.
@@
Given that $\left\|{\mathbf{n}}\right\|^{2}\geq\frac{{\nu}_{0}^{2}+{\nu}_{\max}^{2}}{2}d$ and $\left\|{{\bm{\mu}}_{i}}\right\|^{2}\leq 0.2d$ for all $i\in[k]$, we have $\frac{P^{(i)}({\mathbf{x}})}{P^{(0)}({\mathbf{x}})}\leq\exp(-0.06d)$ for all $i\in[k]$.
@@
Hence, for $\left\|{\mathbf{n}}\right\|^{2}\geq\frac{{\nu}_{0}^{2}+{\nu}_{\max}^{2}}{2}d$ and $\left\|{{\bm{\mu}}_{i}}\right\|^{2}\leq 0.2d\leq\frac{{\nu}_{0}^{2}-{\nu}_{i}^{2}}{2}\left(\log\left(\frac{{\nu}_{i}^{2}}{{\nu}_{0}^{2}}\right)-\frac{{\nu}_{i}^{2}}{2{\nu}_{0}^{2}}+\frac{{\nu}_{0}^{2}}{2{\nu}_{i}^{2}}\right)d$, we have
@@
$\leq\exp\left(-\left(\log\left(\frac{{\nu}_{i}^{2}}{{\nu}_{0}^{2}}\right)-\frac{{\nu}_{i}^{2}}{2{\nu}_{0}^{2}}+\frac{{\nu}_{0}^{2}}{2{\nu}_{i}^{2}}\right)\frac{d}{4}\right)\leq\exp(-0.06d).$
Therefore we finish the proof of Lemma 10. ∎
@@
Suppose $\delta_{t}\leq{\nu}_{0}^{2}$ and $\left\|{\mathbf{n}_{t-1}}\right\|^{2}>36{\nu}_{0}^{2}d$, then for $\mathbf{n}_{t}$ following from equation 13, we have $\left\|{\mathbf{n}_{t}}\right\|^{2}\geq{\nu}_{0}^{2}d$ with probability at least $1-\exp(-0.02d)$.
@@
$=\mathbf{n}_{t-1}-\frac{\delta_{t}}{2}\sum_{i=0}^{k}\frac{w_{i}P^{(i)}({\mathbf{x}}_{t-1})}{P({\mathbf{x}}_{t-1})}\cdot\frac{{\mathbf{N}}^{T}({\mathbf{x}}_{t-1}-{\bm{\mu}}_{i})}{{\nu}_{i}^{2}}+\sqrt{\delta_{t}}\bm{\epsilon}_{t}^{(\mathbf{n})}$
$=\left(1-\frac{\delta_{t}}{2}\sum_{i=0}^{k}\frac{w_{i}P^{(i)}({\mathbf{x}}_{t-1})}{P({\mathbf{x}}_{t-1})}\cdot\frac{1}{{\nu}_{i}^{2}}\right)\mathbf{n}_{t-1}+\sqrt{\delta_{t}}\bm{\epsilon}_{t}^{(\mathbf{n})}.$ (14)
By Lemma 10, we have $\frac{P^{(i)}({\mathbf{x}}_{j-1})}{P^{(0)}({\mathbf{x}}_{j-1})}\leq\exp(-0.06d)$ for all $i\in[k]$, therefore
$1-\frac{\delta_{t}}{2}\sum_{i=0}^{k}\frac{w_{i}P^{(i)}({\mathbf{x}}_{t-1})}{P({\mathbf{x}}_{t-1})}\cdot\frac{1}{{\nu}_{i}^{2}}\geq 1-\frac{\delta_{t}}{2}\cdot\frac{1}{{\nu}_{0}^{2}}-\frac{\delta_{t}(1-w)}{2{\nu}_{i}^{2}w}\exp(-0.06d)>\frac{1}{3}.$ (15)
@@
with probability at least $1-\frac{8}{\sqrt{2\pi d}}\exp\left(-\frac{d}{32}\right)\geq 1-\exp(-0.02d)$. This proves Lemma 11. ∎
@@
For any $t\geq 1$, given that $\delta_{j}\leq{\nu}_{0}^{2}$ and $\frac{{\nu}_{0}^{2}+{\nu}_{\max}^{2}}{2}d\leq\left\|{\mathbf{n}_{j-1}}\right\|^{2}\leq 36{\nu}_{0}^{2}d$ for all $j\in[t]$ and $\left\|{{\bm{\mu}}_{i}}\right\|^{2}\leq 0.2d$ for all $i\in[k]$, we have $\left\|{\hat{\mathbf{n}}_{t}-\mathbf{n}_{t}}\right\|\leq\frac{t}{\exp(0.04d)}\sqrt{d}$.
@@
$=\left\|{\left(1-\frac{\delta_{j}}{2{\nu}_{0}^{2}}\right)(\hat{\mathbf{n}}_{j-1}-\mathbf{n}_{j-1})+\frac{\delta_{j}}{2}\sum_{i\in[k]}\frac{w_{i}P^{(i)}({\mathbf{x}}_{j-1})}{P({\mathbf{x}}_{j-1})}\left(\frac{1}{{\nu}_{i}^{2}}-\frac{1}{{\nu}_{0}^{2}}\right)\mathbf{n}_{j-1}}\right\|$
$\leq\left(1-\frac{\delta_{j}}{2{\nu}_{0}^{2}}\right)\left\|{\hat{\mathbf{n}}_{j-1}-\mathbf{n}_{j-1}}\right\|+\sum_{i\in[k]}\frac{\delta_{j}}{2}\frac{w_{i}P^{(i)}({\mathbf{x}}_{j-1})}{P({\mathbf{x}}_{j-1})}\left(\frac{1}{{\nu}_{i}^{2}}-\frac{1}{{\nu}_{0}^{2}}\right)\left\|{\mathbf{n}_{j-1}}\right\|$
$\leq\left\|{\hat{\mathbf{n}}_{j-1}-\mathbf{n}_{j-1}}\right\|+\sum_{i\in[k]}\frac{\delta_{j}}{2}\frac{w_{i}P^{(i)}({\mathbf{x}}_{j-1})}{w_{0}P^{(0)}({\mathbf{x}}_{j-1})}\left(\frac{1}{{\nu}_{i}^{2}}-\frac{1}{{\nu}_{0}^{2}}\right)6{\nu}_{0}\sqrt{d}.$
By Lemma 10, we have $\frac{P^{(i)}({\mathbf{x}}_{j-1})}{P^{(0)}({\mathbf{x}}_{j-1})}\leq\exp(-0.06d)$ for all $i\in[k]$, hence we obtain a recursive bound
$\left\|{\hat{\mathbf{n}}_{j}-\mathbf{n}_{j}}\right\|\leq\left\|{\hat{\mathbf{n}}_{j-1}-\mathbf{n}_{j-1}}\right\|+\frac{1}{\exp(0.04d)}\sqrt{d}.$
@@
$\left\|{\hat{\mathbf{n}}_{t}-\mathbf{n}_{t}}\right\|=\sum_{j\in[t]}\left(\left\|{\hat{\mathbf{n}}_{j}-\mathbf{n}_{j}}\right\|-\left\|{\hat{\mathbf{n}}_{j-1}-\mathbf{n}_{j-1}}\right\|\right)\leq\frac{t}{\exp(0.04d)}\sqrt{d}.$
@@
Given that $\left\|{\hat{\mathbf{n}}_{0}}\right\|^{2}\geq\frac{3{\nu}_{0}^{2}+{\nu}_{\max}^{2}}{4}d$, we have $\left\|{\hat{\mathbf{n}}_{t}}\right\|^{2}\geq\frac{5{\nu}_{0}^{2}+3{\nu}_{\max}^{2}}{8}d$ with probability at least $1-\exp\left(-d/300\right)$.
@@
$\leq{\mathbb{P}}\left(\left\|{\bm{\epsilon}}\right\|^{2}\leq n-2\sqrt{n\left(\frac{{\nu}_{0}^{2}-{\nu}_{\max}^{2}}{8{\nu}_{0}^{2}}\right)^{2}\frac{d}{2}}\right)$
$\leq\exp\left(-\left(\frac{{\nu}_{0}^{2}-{\nu}_{\max}^{2}}{8{\nu}_{0}^{2}}\right)^{2}\frac{d}{2}\right)\leq\exp(-d/288),$
@@
Upon having all the above lemmas, we are now ready to establish Theorem 2 by induction. Suppose the theorem holds for all $T$ values of $11,\cdots,T-$ 1. We consider the following 3 cases:
  • If there exists some $t\in[T]$ such that $\delta_{t}>{\nu}_{0}^{2}$, by Lemma 9 we know that with probability at least $1-\exp(-d/25)$, we have $\left\|{\mathbf{n}_{t}}\right\|^{2}\geq\frac{3{\nu}_{0}^{2}+{\nu}_{\max}^{2}}{4}d$, thus the problem reduces to the two sub-arrays $\mathbf{n}_{0},\cdots,\mathbf{n}_{t-1}$ and $\mathbf{n}_{t},\cdots,\mathbf{n}_{T}$, which can be solved by induction.
  • Suppose $\delta_{t}\leq{\nu}_{0}^{2}$ for all $t\in[T]$. If there exists some $t\in[T]$ such that $\left\|{\mathbf{n}_{t-1}}\right\|^{2}>36{\nu}_{0}^{2}d$, by Lemma 11 we know that with probability at least $1-\exp(-d/50)$, we have $\left\|{\mathbf{n}_{t}}\right\|^{2}\geq{\nu}_{0}^{2}d>\frac{3{\nu}_{0}^{2}+{\nu}_{\max}^{2}}{4}d$, thus the problem similarly reduces to the two sub-arrays $\mathbf{n}_{0},\cdots,\mathbf{n}_{t-1}$ and $\mathbf{n}_{t},\cdots,\mathbf{n}_{T}$, which can be solved by induction.
@@
    By Lemma 14 we have that with probability at least $1-\exp(-d/300)$,